A railroad curve is laid out on a circle. What radius would be used if the track is to change direction by 20 degrees in a distance of 100 km?
I've been trying to use formula r=s/theta but I can't find the solution for theta.
Please help! Thanks.

Answer by kapilsinghi(68)

if see it as and arc of circle whose radius is r Km
Circumfrence of this circle is 2*pi*r
to turn 20 Degrees it has to cover 100Km
to turn 360 Degrees it has to cover 360*100/20 = circumfrence of the circle
360*5 = 2*3.14*r
1800/6.28 = r
286.64 = r
radius of this circle is 286.64 KM

kapilsinghi123@gmail.com